OpenStudy (anonymous):

Rationalize the denominator. Assume that all expressions under radicals represent positive numbers. (√r-√s)/(√r+√s)

OpenStudy (anonymous):

You will want to multiple by the fraction (√r - √s)/(√r - √s)

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt{r}-\sqrt{s}\]

OpenStudy (anonymous):

denominator will be \[r-s\] numerator will be whatever you get when you multiply out, namely \[r+s-2\sqrt{rs}\]

OpenStudy (anonymous):

the bottom will simplify to just being r-s. the numerator should be (r + s - 2√(rs)

OpenStudy (anonymous):

what TBates said